You roll two dice and total the up faces. What is the probability of getting a total of 8 or two up faces that are the same?
The answer is 0.28 (28%) but i would love to understand why this is so.

Answer by scott8148(6628)

there are 36 ways for the dice to land
___ each of 6 faces on one die combined with the 6 faces on the other die ___ 6 * 6 = 36

to get 8:
2,6
3,5
4,4
5,3
6,2
___5 ways

6 ways to get the faces the same (numbers one through six)

the 4,4 is in both groups but only counts once ___ so there are 10 ways to get an 8 or matching numbers

10 out of 36 = 10/36 = .278 (approx)